# A Compact Low-level RF Control System for Advanced Concept Compact Electron Linear Accelerator

### v1

@@ Abstract
A compact low-level RF (LLRF) control system based on RF system-on-chip (RFSoC) technology has been designed for the Advanced Concept Compact Electron Linear-accelerator (ACCEL) program, which has challenging requirements in both RF performance and size, weight and power consumption (SWaP). The compact LLRF solution employs the direct RF sampling technique of RFSoC, which samples the RF signals directly without any analogue up and down conversion. Compared with the conventional heterodyne based architecture used for LLRF system of linear accelerator (LINAC), the elimination of analogue mixers can significantly reduce the size and weight of the system, especially with LINAC requires a larger number of RF channels. Based on the requirements of ACCEL, a prototype LLRF platform has been developed, and the control schemes have been proposed. The prototype LLRF system demonstrated magnitude and phase fluctuation levels below 1% and 1° on the flat top of a 2 $\mu$s RF pulse. The LLRF control schemes proposed for ACCEL are implemented with a prototype hardware platform. This paper will introduce the new compact LLRF solution and summarize a selection of experimental test results of the prototype itself and with the accelerating structure cavities designed for ACCEL.
@@ I Introduction
The Advanced Concept Compact Electron Linear-accelerator (ACCEL) program aims to develop a high-power compact, rugged linear accelerator that can be easily transported and deployed in the field. The most critical and challenging requirement of the ACCEL is size, weigh and power consumption (SWaP) of the linear accelerator (LINAC) DARPA . All the RF components of ACCEL, including the RF source, amplification, acceleration structure and low-level RF (LLRF), have been designed to achieve high RF power with stringent SWaP requirements. In this paper, the prototype compact LLRF system for ACCEL, including the hardware platform and control schemes, will be introduced and discussed.
The ACCEL system has been designed to operate in C-band around 5.712 GHz. The accelerating structure has 26 pairs of cavities and with each pairing sharing a single RF input and two RF monitors. With conventional LLRF system with heterodyne based architecture, at least 78 analogue up and down mixing circuits with discrete data converters would be required. The size and weight of the LLRF hardware will be substantial after including the field programmable gate array (FPGA) and processing system (PS) for control and interfacing. The compact LLRF system for ACCEL has been designed based on the RF system-on-chip (RFSoC) device family from AMD Xilinx. The RFSoC device integrates RF data converters, FPGA and PS, which are the essential components of a LLRF system. The integrated data converters in RFSoC devices offer higher order Nyquist zone sampling, which enable the C-band RF signals to be sampled or generated directly without analogue mixing. The RFSoC device selected for the compact LLRF, AMD Xilinx Zynq UltraScale+ RFSoC Gen 3 ZU49DR, has 16 RF analogue to digital converters (ADCs) with maximum RF input frequency of 6 GHz and a maximum sampling rate of 2.5 GHz and 16 RF digital to analogue data converters (DACs) with a maximum sampling rate of 9.85 GHz AMD . The LLRF system for the complete ACCEL structure can be realized with 4 RFSoC devices with minimum other electronics. The high integration level of the RFSoCs offers substantial benefits in the overall SWaP of ACCEL.
@@
The RFSoC technology has been widely used in research and development of control and readout platforms for a range of physics experiments at SLAC National Accelerator Laboratory and other institutes we collaborate with Liu, Jones, and Taylor (2021, 2022); Liu, Ruckman, and Herbst (2023); Henderson et al. (2022); Liu et al. (2023, 2024a). The LLRF for the cool cooper collider (C3), which also operates in C-band, has been prototyped based on an RFSoC device in the same family as part of the road map for C3 R&D Bai et al. (2021); Vernieri et al. (2023); Nanni et al. (2023). The initial RF performance evaluation of the LLRF for C3 with a solid state amplifier (SSA) demonstrated magnitude and phase fluctuation levels within 1 s are 0.34 % and 0.37° respectively Liu et al. (2024b), which is considerably better than the 1% and 1°  requirements for ACCEL. In Liu et al. (2024c), the next generation LLRF (NG-LLRF) platform based on RFSoC developed at SLAC has been introduced. The NG-LLRF platform demonstrated that it can drive and measure a high-power test setup with high precision. The compact LLRF system for ACCEL has been designed and implemented based on configuration optimization and performance characterization exercises performed for RFSoCs with the other LLRF systems of LINACs and other physics experiments. In Section II, the architecture of the prototype compact LLRF system, including hardware, firmware and software, will be introduced.
@@
An LLRF system typically employs a pulse-to-pulse feed-back loop to stabilize the RF field in an accelerating structure Geng (2017). The pulse-to-pulse magnitude and phase control scheme proposed will be described in section III.3. Due to the special operating conditions of ACCEL, more control schemes are required. One of the design requirements of ACCEL is to operate in a wide temperature range without temperature stabilization for the accelerating structure. The resonant frequency of the accelerating structure can drift with the temperature substantially, which can affect the RF power injection to the structure. We have designed an RF frequency tuning scheme to track the resonance of structure cavities and change the RF frequency accordingly. The RF frequency tuning scheme will be described in Section III.1. The ACCEL system is designed to be operated with an electron bunch train within an RF pulse, which require the phase and magnitude to be stable within each pulse. The scheme proposed to control magnitude and phase within the pulse will be described in Section III.2.
@@
The compact LLRF prototype has been tested at some critical stages of the development. The first test with the prototype has been performed is the direct loopback test with platform and the fluctuation levels of the loopback setup define the highest stability level of entire control system. The test results of loopack test setup will be summarized in Section IV.1. The RF frequency tuning scheme has been implemented and tested with the ACCEL structure cavities. The RF signal before and after the RF frequency has been tuned to the resonance of the accelerating structure will be demonstrated and discussed in Section IV.2. A prototype section of the compact LLRF system has also been tested will the high-power C-band test facility with a water load and a selection of the test results will be summarized in Section IV.3.
@@ II Compact LLRF Control Hardware Platform
The prototype compact LLRF system for ACCEL has been developed based on the ZCU216 evaluation board, which carries an RFSoC Gen 3 ZU49DR device. A block diagram for the platform is shown in Figure 1. The solid line in Figure 1 defines the boundary between the RFSoC and the external components and signals. Within the RFSoC, the system implementation has been divided to three parts separate by the dashed lines: the RF data converter (RFdc), the programmable logic (PL) and the processing system (PS). The mixers and decimation or interpolation blocks before or after the data converter are integrated as a part of RFdc and each of the chain is named as a ‘datapath’. The datapaths for the data converters can be configured independently from each other.
@@
For each pair of the ACCEL cavities, there are two RF signals to be measured and used for control and monitoring purposes, which are the klystron forward signal and cavity reflection signal. As Figure 1 shows, the RF signals are filtered with a band-pass filter and then sampled directly by the integrated ADCs in RFSoC at 2.4576 giga samples per second (GSPS). The digitized signal is then down converted to 798.6 MHz, which is the image of 5.712 GHz in the first Nyquist zone. The base-band signal in in-phase (I) and quadrature (Q) components are decimated by a factor of 10. Therefore, the bandwidth of the RF measurement chain in this case is 245.76 MHz. The bandwidth can be adjusted for future requirement by using a different decimation factor. The device uses digital finite impulse response (FIR) filters to reject high frequency noise as a part of the decimation process.
@@
For the ACCEL structure design, there are no probes in the cavity structure. The cavity reflection is the only signal that carries information for the field inside the structure. In this version of the prototype, the IQ components of the cavity reflection signal in streaming format has converted to memory-mapped format and written to the DDR, which is outside of RFSoC chip. Data can then be accessed by the PS and sent to a server or processed locally on the PS. For the experiments in this paper, the data are sent to a server via Ethernet for further processing.
@@
The down-converted klystron forward signal in IQ format is streamed to the feedback control algorithm firmware block. The user input parameters from the software layer are loaded to the firmware block via registers. Based on the IQ components and user input parameters, the control block calculates the updated IQ pair that compensate for the pulse-to-pulse fluctuation in phase and magnitude levels. The user can load a custom waveform via software, otherwise the default square wave will be used. The waveform is modulated with the updated IQ in the modulation block, which generates the updated base-band pulse. The DAC datapath interpolates the base-band pulse first and then up mixes with the digital values from the numerical controlled oscillator (NCO). The digital up conversion frequency can also be controlled via software. The DAC decodes and generates the updated RF pulse with a sample rate of 5.89824 GHz and drives the entire structure.
@@ III LLRF Control Schemes
The ACCEL LLRF control system has been divided into three main parts based on the requirements. The control schemes proposed are designed to be performed in order: RF frequency control, followed by pulse shape control, and then the pulse-to-pulse phase and magnitude control feedback loop. The control schemes proposed are demonstrated in the form of flow charts in the following three subsections. In the flow charts, the orange blocks are planned be implemented in software and blue blocks to be implemented in firmware. The interface between the software and firmware will be defined in the Section III.4.
@@ III.1 RF frequency tuning
The targeted operation temperature range of ACCEL is from -40 to 85 °C. The resonant frequency of the accelerating changes by around 0.1 MHz per °C. Due to the limits in size and weight, there is no temperature stabilization mechanism for the accelerating structure of ACCEL and maintaining beam and beam energy is the priority of the program. Therefore, the RF frequency needs to be tuned according to the resonance frequency of the structure while operating.
@@
Figure 2 shows the flow chart of RF frequency control flow. When the system is first switched on, the RF frequency control flow will be executed. The RF frequency will be coarsely set based on the reading of the temperature sensor and the look up table of temperature verse RF frequency that is embedded in the software. With the coarse RF frequency tuning, a fraction of RF power should be able to be injected into the cavity and the energy stored in the structure will radiate at the resonant frequency of the structure after the RF pulse is off. The radiation can be captured by the cavity reflection coupler and the IQ samples of the cavity reflection signal from the ADC datapath is sent to software. Then the frequency error can be calculated based on the phase of the cavity reflection signal after the RF pulse is off and the calculation theory and experimental results will be elaborated in section IV.2. The updated RF frequency calculated by software is loaded to the data converter datapaths and changes the frequency of the sequence generated by the NCOs, which will ultimately update the RF frequency for both the ADCs and DACs. The process frequency correction can be executed in a controlled step size until the frequency error is below a desired value defined by the user. Based on the gradient of temperature change while operation, the frequency tuning loop can be executed at an adaptive rate.
@@ III.2 RF pulse shape control
Figure 3 shows the flow chart of the pulse shape control scheme. After the RF frequency is set to the resonance of the accelerating structure by the RF frequency tuning loop, the pulse shape control flow will be executed. As there is no probe that measures the RF field in the structure directly and the RF field follows the klystron by design, the klystron forward signal is employed for RF pulse shape control. The magnitude and phase values of the klystron forward signal are sent to software to calculate the updated pulse shape that can compensate for any fluctuation within the pulse. In software, the updated pulse shape is calculated based on the variation between the desired and measured pulse shape to reduce the offset between them sample by sample. Then the updated base-band pulse shape in IQ format is loaded to the firmware and upconverted to drive the system for the following pulse. The procedure repeats until the standard deviations of magnitude and phase on flat top are below the desired levels. The pulse shape should be consistent after the compensation procedure is performed, so the pulse shape control may not required to executed continuously after the fist run. However, more development and test efforts are required to verify the optimal control routine and algorithm.
@@ III.3 Pulse-to-pulse Magnitude and Phase Control
Measuring and stabilizing the magnitude and phase of the field in the accelerating structure is the primary objective for most of LLRF control systems of LINACs. Figure 4 shows the flow chart of phase and magnitude control we have proposed for ACCEL. After the RF frequency and pulse-top flatness are tuned to desire levels, the phase and magnitude feedback control loops will start operating. This control loop is also be implemented by using the klystron forward signal. In firmware, the base-band klystron forward signal from ADC datapath will be processed with moving average with certain number of samples to improve the signal stability. The IQ components are converted to phase and magnitude values and the average values will be computed. The current average magnitude and phase values will be compared to the user defined values and then calculating for a new set of magnitude and phase values to approach the user defined values in the defined rates. The new set of magnitude and phase values are converted back to IQ format and modulated with the base-band pulse waveform. The up converted RF signal samples are finally clocked into the DAC to generate the updated RF pulse, which will be used as the drive of accelerating structure for the following pulse. The feedback loop runs continuously during operation to perform real-time pulse to pulse magnitude and phase correction. If the difference between the measurement and desired values are below a certain threshold, the values will not be changed for the following pulse. This is to minimize the instability introduced by the LLRF if the desired signal levels are already meeting operating requirements.

### v2

@@ Abstract
A compact low-level RF (LLRF) control system based on RF system-on-chip (RFSoC) technology has been designed for the Advanced Concept Compact Electron Linear-accelerator (ACCEL) program, which has challenging requirements in both RF performance and size, weight and power consumption (SWaP). The compact LLRF solution employs the direct RF sampling technique of RFSoC, which samples the RF signals directly without any analogue up and down conversion. Compared with the conventional heterodyne based architecture used for LLRF system of linear accelerator (LINAC), the elimination of analogue mixers can significantly reduce the size and weight of the system, especially with LINAC requires a larger number of RF channels. Based on the requirements of ACCEL, a prototype LLRF platform has been developed and the control schemes have been proposed. The prototype LLRF system demonstrated magnitude and phase fluctuation levels below 1% and 1° on the flat top of a 2 $\mu$s RF pulse. The LLRF control schemes proposed for ACCEL are implemented with a prototype hardware platform. This paper will introduce the new compact LLRF solution and summarize a selection of experimental test results of the prototype itself and with the accelerating structure cavities designed for ACCEL.
@@ I Introduction
The Advanced Concept Compact Electron Linear-accelerator (ACCEL) program aims to develop a high-power compact, rugged linear accelerator that can be easily transported and deployed in the field. The most critical and challenging requirement of the ACCEL is the size, weight and power consumption (SWaP) of the linear accelerator (LINAC) DARPA . All the RF components of ACCEL, including the RF source, amplification, acceleration structure, and low-level RF (LLRF), have been designed to achieve high RF power with stringent SWaP requirements. In this paper, the prototype compact LLRF system for ACCEL, including the hardware platform and control schemes, will be introduced and discussed.
The ACCEL system has been designed to operate in the C-band around 5.712 GHz. The accelerating structure has 26 pairs of cavities, and with each pair shares a single RF input and two RF monitors. With the conventional LLRF system with heterodyne based architecture, at least 78 analogue up and down mixing circuits with discrete data converters would be required. The size and weight of the LLRF hardware will be substantial after including the field programmable gate array (FPGA) and processing system (PS) for control and interfacing. The compact LLRF system for ACCEL has been designed based on the RF system-on-chip (RFSoC) device family from AMD Xilinx. The RFSoC device integrates RF data converters, FPGA and PS, which are the essential components of an LLRF system. The integrated data converters in RFSoC devices offer higher-order Nyquist zone sampling, enabling C-band RF signals to be sampled or generated directly without analog mixing. The RFSoC device selected for compact LLRF, AMD Xilinx Zynq UltraScale+ RFSoC Gen 3 ZU49DR, has 16 RF analogue to digital converters (ADCs) with maximum RF input frequency of 6 GHz and a maximum sampling rate of 2.5 GHz and 16 RF digital to analogue data converters (DACs) with a maximum sampling rate of 9.85 GHz AMD . The LLRF system for the complete ACCEL structure can be realized with 4 RFSoC devices with a minimum of other electronics. The high integration level of the RFSoCs offers substantial benefits in the overall SWaP of the ACCEL.
@@
RFSoC technology has been widely used in the research and development of control and readout platforms for a variety of physics experiments at the SLAC National Accelerator Laboratory and other institutes with which we collaborate Liu, Jones, and Taylor (2021, 2022); Liu, Ruckman, and Herbst (2023); Henderson et al. (2022); Liu et al. (2023, 2024a). The LLRF for the cool cooper collider (C3), which also operates in the C-band, has been prototyped based on an RFSoC device in the same family as part of the road map for C3 R&D Bai et al. (2021); Vernieri et al. (2023); Nanni et al. (2023); Andorf et al. (2025). The initial RF performance evaluation of the LLRF for C3 with a solid state amplifier (SSA) demonstrated that the magnitude and phase fluctuation levels within 1 s are 0.34 % and 0.37° respectively Liu et al. (2024b), which is considerably better than the requirements of 1% and 1° for ACCEL. In Liu et al. (2024c), the next generation LLRF (NG-LLRF) platform based on RFSoC developed at SLAC has been introduced. The NG-LLRF platform demonstrated that it can drive and measure a high-power test setup with high precision in both the C-band Liu et al. (2024b, 2025a) and the S-band Liu et al. (2025b). The compact LLRF system for ACCEL has been designed and implemented based on configuration optimization and performance characterization exercises performed for RFSoCs with the other LLRF systems of LINACs and other physics experiments. In Section II, the architecture of the prototype compact LLRF system, including hardware, firmware, and software, will be introduced.
@@
An LLRF system typically employs a pulse-to-pulse feedback loop to stabilize the RF field in an accelerating structure Geng (2017). The proposed pulse-to-pulse magnitude and phase control scheme will be described in Section III.3. Due to the special operating conditions of ACCEL, more control schemes are required. One of the design requirements of the ACCEL is to operate in a wide temperature range without temperature stabilization for the accelerating structure. The resonant frequency of the accelerating structure can drift with temperature, which can affect the injection of RF power into the structure. We have designed an RF frequency tuning scheme to track the resonance of the structure cavities and change the RF frequency accordingly. The RF frequency tuning scheme will be described in Section III.1. The ACCEL system is designed to operate with an electron bunch train within an RF pulse, requiring the phase and magnitude to be stable within each pulse. The scheme proposed to control the magnitude and phase within the pulse will be described in Section III.2.
@@
The compact LLRF prototype has been tested in some critical stages of t development. The first test with the prototype has been performed is the direct loopback test with platform and the fluctuation levels of the loopback setup define the highest stability level of entire control system. The test results of the loopack test setup will be summarized in Section IV.1. The RF frequency tuning scheme has been implemented and tested with the ACCEL structure cavities. The phase and magnitude of the cavity reflection signals before and after the tuning scheme applied will be demonstrated and discussed in Section IV.2. A prototype of the compact LLRF system has also been tested at the high-power C-band test facility with a water load. A selection of test results will be summarized in Section IV.3.
@@ II Compact LLRF Control Hardware Platform
The compact LLRF prototype for ACCEL has been developed based on the ZCU216 evaluation board, which carries an RFSoC Gen 3 ZU49DR device. A block diagram for the platform is shown in Figure 1. The solid line in Figure 1 defines the boundary between the RFSoC and the external components and signals. Within the RFSoC, the system implementation has been divided into three parts separate by the dashed lines: the RF data converter (RFdc), the programmable logic (PL) and the processing system (PS). The mixers and decimation or interpolation blocks before or after the data converter are integrated as a part of RFdc and each of the chains is named as a ‘datapath’. The datapaths for the data converters can be configured independently of each other.
@@
For each pair of the ACCEL cavities, there are two RF signals to be measured and used for control and monitoring purposes, which are the klystron forward signal and the cavity reflection signal. As Figure 1 shows, RF signals are filtered with a band-pass filter and then sampled directly by the integrated ADCs in RFSoC at 2.4576 giga samples per second (GSPS). The digitized signal is then down converted to 798.6 MHz, which is the image of 5.712 GHz in the first Nyquist zone. The base-band signal in in-phase (I) and quadrature (Q) format is decimated by a factor of 10. Since the signal is quadrature sampled after the digital down-mixing, the bandwidth of the RF measurement chain in this case is 245.76 MHz, which is the same as the data rate. The bandwidth can be adjusted for future requirements using a different decimation factor. The device uses digital finite impulse response (FIR) filters to reject high frequency noise as a part of the decimation process.
@@
For the ACCEL structure design, there are no probes in the cavity structure. The cavity reflection is the only signal that carries information about the field inside the structure. In this version of the prototype, the IQ components of the cavity reflection signal in streaming format have been converted to memory-mapped format and written to the DDR, which is outside of the RFSoC. The data can then be accessed by the PS and sent to a server or processed locally on the PS. For the experiments in this paper, the data was sent to a server via Ethernet for further processing.
@@
The down-converted klystron forward signal in IQ format is streamed to the feedback control algorithm firmware block. The user input parameters from the software layer are loaded to the firmware block via registers. Based on the IQ components and user input parameters, the control block calculates the updated IQ pair that compensates for the pulse-to-pulse fluctuation in phase and magnitude levels. The user can load a custom waveform via software; otherwise the default square wave will be used. The waveform is modulated with the updated IQ in the modulation block, which generates the updated base-band pulse. The DAC datapath interpolates the base-band pulse first and then up mixed with the digital values from the numerical controlled oscillator (NCO). The digital up conversion frequency can also be controlled via software. The DAC decodes and generates the updated RF pulse with a sample rate of 5.89824 GHz and drives the entire structure.
For the ZCU216 platform we selected, the master clock for the RFSoC is generated by an external phase locked loop (PLL) device from a reference signal. In this case, we have only one RFSoC, so an on-board 10 MHz oscillator is used as a reference. However, when there are multiple RFSoCs in the system, all devices need clocks that are frequency locked and phase aligned. For each of the RFSoCs, the master clock is delivered to one of the data converter tiles and propagated to other tiles using an internal network. The clocks for sampling, NCO and filtering blocks of each tile are generated by internal PLLs from the master clock propagated to the tile. Therefore, all integrated blocks of the RFSoC devices are synchronized with the master clock. The propagation delays between the data converters and the programming logic can be handled by the Multiple Tile Synchronization (MTS) technique of the RFSoC family. With MTS, the latency between tiles can be deterministically aligned between reset and reset. Reset of NCOs from different tiles or different devices should be triggered by a common system reference to achieve repeatable operation.
@@ III LLRF Control Schemes
The ACCEL LLRF control system has been divided into three main parts based on the requirements. The control schemes proposed are designed to be performed in order: RF frequency control, followed by pulse shape control, and then the pulse-to-pulse phase and magnitude control feedback loop. The control schemes proposed are demonstrated in the form of flow charts in the following three subsections. In the flow charts, the orange blocks are planned to be implemented in software and the blue blocks to be implemented in firmware. The interface between the software and the firmware will be defined in Section III.4.
@@ III.1 RF frequency tuning
The target operating temperature range of ACCEL is from -40 to 85 °C. The resonant frequency of the accelerating changes by around 0.1 MHz per °C. Due to the limits in size and weight, there is no temperature stabilization mechanism for the accelerating structure of ACCEL and maintaining beam and beam energy is the priority of the program. Therefore, the RF frequency needs to be tuned according to the resonance frequency of the structure while operating.
@@
Figure 2 shows the flow chart of the RF frequency control flow. When the system is first switched on, the RF frequency tuning flow will be executed. The RF frequency will be coarsely set based on the reading of the temperature sensor and the lookup table of the temperature versus RF frequency embedded in the software. With the coarse RF frequency tuning, a fraction of RF power should be able to be injected into the cavity, and the energy stored in the structure will radiate at the resonant frequency of the structure after the RF pulse is off. The radiation can be captured by the cavity reflection coupler, and the IQ samples of the cavity reflection signal from the ADC datapath are sent to software. The frequency error can then be calculated based on the phase of the cavity reflection signal after the RF pulse is off, and the calculation theory and experimental results will be elaborated in Section IV.2. The updated RF frequency calculated by the software is loaded into the datapaths of the data converters and changes the frequency of the sequence generated by the NCOs, which will ultimately update the RF frequency for both the ADCs and DACs. The frequency correction process can be executed in a controlled step size until the frequency error is below a desired value defined by the user. The RF frequency tuning will be performed each time the structure is switched on. But it will also be executed when the ambient temperature changes and introduces a detectable structure resonance frequency change above a user-defined threshold. Based on the gradient of temperature change during operation, the frequency tuning loop can be executed at an adaptive rate.
@@ III.2 RF pulse shape control
Figure 3 shows the flow chart of the pulse shape control scheme. After the RF frequency tuning scheme is applied to the system, the pulse shape control flow will be executed. As there is no cavity probe that measures the RF field in the structure directly and the RF field follows the klystron by design, the cavity forward and reflected signals could be used to determine the field strength in the cavity and control the needed shape of the RF pulse. The magnitude and phase values of the klystron forward signal are sent to the software to calculate the updated pulse shape that can compensate for any fluctuation within the pulse. In software, the updated pulse shape is calculated based on the variation between the desired and measured pulse shape to reduce the offset between them sample by sample. Then the updated base-band pulse shape in IQ format is loaded to the firmware and up-converted to drive the system for the following pulse. The procedure repeats until the standard deviations of magnitude and phase on the flat top are below the desired levels. The pulse shape should be consistent after the compensation procedure is performed, so the pulse shape control may not require being executed continuously after the first run. In Liu et al. (2025a, c), we summarize the test results of different RF pulse modulation schemes with an LLRF system based on RFSoC integrated with high-power test stands in the C-band. The system demonstrates high precision in the generation and measurement of RF pulses with amplitude and phase modulation schemes. However, operating the accelerator with an arbitrary waveform is the first step of pulse shaping, and more development and test efforts are required for the control routine and algorithm to deliver the field with optimum stability to the accelerating structure. Since the pulse shape may change with the drifts of the RF components, the stability of the control scheme should be evaluated in the next phase of development.
@@ III.3 Pulse-to-pulse Magnitude and Phase Control
Measurement and stabilization of the magnitude and phase of the field in the accelerating structure are the primary objectives for most of the LLRF control systems of LINACs. Figure 4 shows the flow chart of phase and magnitude control that we have proposed for ACCEL. After the RF frequency and pulse-top flatness are tuned to desired levels, the phase and magnitude feedback control loops will start operating. This control loop is also be implemented by using the klystron forward signal. In firmware, the base-band klystron forward signal from the ADC datapath will be processed with moving average with a certain number of samples to improve the signal stability. The IQ components are converted to phase and magnitude values, and the average values will be computed. The current average magnitude and phase values will be compared to the user defined values and then calculating for a new set of magnitude and phase values to approach the user defined values in the defined rates. The new set of magnitude and phase values is converted back to IQ format and modulated with the base-band pulse waveform. The up converted RF signal samples are finally clocked into the DAC to generate the updated RF pulse, which will be used as the drive of the accelerating structure for the following pulse. The feedback loop runs continuously during operation to perform real-time pulse to pulse magnitude and phase correction. If the difference between the measurement and the desired value is below a certain threshold, the values will not be changed for the following pulse. This is done to minimize the instability introduced by the LLRF if the desired signal levels are already meeting the operating requirements.